### v1

# Inductive van der Waals Force between Two Quantum Loops

Abstract
We study the van der Waals-London force, which is typically associated with fluctuating dipoles in atoms, in a mesoscopic circuit consisting of two inductively coupled superconducting loops. We investigate the inductive van der Waals-London interaction using both semiclassical and quantum electrodynamic (QED) approaches. The semiclassical model predicts a repulsive interaction due to anticorrelated current fluctuations. In contrast, the QED framework, which incorporates virtual photon exchange, reveals a predominantly attractive force. A key contribution comes from a state-independent two-photon exchange, which is absent in the semiclassical description and undetectable by spectroscopy. Our study introduces a new experimental platform for measuring the van der Waals force between individual artificial atoms via controlled mesoscopic circuits.
Introduction-. The van der Waals-London interaction between two neutral atoms or molecules is ubiquitous in nature and underlies a wide variety of physical phenomena Parsegian (2005). Microscopically, it results from the interaction of fluctuating electric dipoles London (1937). The van der Waals (vdW) force is macroscopic in manifestation. However, measuring it between individual atoms remains unachieved. Although indirect spectroscopic measurements have been achieved between two Rydberg atoms Beguin et al. (2013), directly detecting the force is more challenging.
We note that mesoscopic electronic circuits are an ideal platform for studying van der Waals forces between individual “atoms”, since artificial atoms can be fabricated on mesoscopic or even macroscopic scales. Their larger dipole moments lead to significantly stronger interatomic interactions (see, for example, Ref. You and Nori, 2011). Due to remarkable progress in sensitive measurement technology (see, for example, Ref. Degen et al., 2017), we expect the vdW force between individual artificial atoms to be directly observable. Additionally, electric dipole interactions dominate in real atoms, but artificial atoms can be designed to have dominant magnetic (inductive) interactions.
In this Letter, we study the vdW-London interaction in a system of two inductively coupled quantum loops (see Fig. 1), which can be realized with superconducting condensates. Classically, the inductive (magnetic) force between the loops is proportional to the currents, and thus, it vanishes in the absence of current flow. This system enables the study of the van der Waals-London interaction, which arises from quantum fluctuations in the current of the zero-current ground state. We find that the behavior of the inductive vdW-London interaction differs from the standard electric dipole-dominated case. Using a quantum electrodynamic(QED) approach, we demonstrate that inductive vdW-London interactions exhibit more complex properties, primarily due to the two-photon process of the “atom”-photon interaction Hamiltonian (the $\mathbf{A}^{2}$ term, where $\mathbf{A}$ is the vector potential of the radiation field). An interesting geometric feature of the Fig. 1 setup is that the distance $R$ between the loops can be much shorter than their size. This condition is unattainable with real atoms in three dimensions.
Figure 1: Schematic diagrams of the two interacting quantum loops: (a) In the semiclassical model, the loops are represented by an inductive coupling of the two quantum LC circuits. (b) In QED, vacuum photons mediate the interaction between the loops.
Semiclassical approach with LC circuit model-. Two inductively coupled quantum loops (Fig. 1(a)) can be modeled by the Lagrangian
${\cal L}={\cal L}_{0}+{\cal L}_{i}\,$, (1a)
where
${\cal L}_{0}=\frac{L}{2}\dot{Q}_{1}^{2}-\frac{Q_{1}^{2}}{2C}+\frac{L}{2}\dot{Q}_{2}^{2}-\frac{Q_{2}^{2}}{2C}$ (1b)
and
${\cal L}_{i}=M\dot{Q}_{1}\dot{Q}_{2}$ (1c)
represent each loop and the mutual interaction, respectively. Here, $Q_{i}$ denotes the charge at circuit $i$ ($i=1,2$), and its time derivative, $\dot{Q}_{i}$, is the current. For simplicity, the two loops are considered identical, each with inductance $L$ and capacitance $C$, and the angular frequency is $\omega_{0}=1/\sqrt{LC}$. The mutual inductance,
$M(\mathbf{R})=\frac{\mu_{0}}{4\pi}\oint_{1}\oint_{2}\frac{d\mathbf{r}_{1}\cdot d\mathbf{r}_{2}}{|\mathbf{r}_{1}-\mathbf{r}_{2}|},$ (2)
depends on the loop geometry and the displacement vector $\mathbf{R}$. The Euler-Lagrange equation yields the mutual force:
$\mathbf{F}(\mathbf{R})=\frac{\partial\cal L}{\partial\mathbf{R}}=\frac{\partial M}{\partial\mathbf{R}}\dot{Q}_{1}\dot{Q}_{2}\,.$ (3)
This classical force between the loops disappears in the absence of current flow. However, in the quantum regime, with macroscopic condensates in each loop, quantum fluctuations of the current play an essential role. Using the transformation
$Q_{\pm}=\frac{1}{\sqrt{2}}(Q_{1}\pm Q_{2}),$ (4a)
the Lagrangian (Eq. (1)) becomes
${\cal L}=\frac{L+M}{2}\dot{Q}_{+}^{2}-\frac{Q_{+}^{2}}{2C}+\frac{L-M}{2}\dot{Q}_{-}^{2}-\frac{Q_{-}^{2}}{2C}\,.$ (4b)
The corresponding Hamiltonian is
$H=\frac{L+M}{2}I_{+}^{2}+\frac{Q_{+}^{2}}{2C}+\frac{L-M}{2}I_{-}^{2}+\frac{Q_{-}^{2}}{2C}\,.$ (5)
where $I_{\pm}=\frac{1}{L\pm M}\frac{\partial\cal L}{\partial\dot{Q}_{\pm}}$ is the current variable associated with $Q_{\pm}$. This represents two independent oscillators of the variables $Q_{+}$ and $Q_{-}$ with the angular frequencies
$\omega_{\pm}=\omega_{0}(1\pm\eta)^{-1/2},$ (6)
where $\eta\equiv M/L$.
The ground state energy from this semiclassical treatment ($E_{\mathrm{sc}}$) is the sum of the zero-point energies:
$E_{\mathrm{sc}}=\frac{\hbar}{2}\left(\omega_{+}+\omega_{-}\right)=\frac{\hbar\omega_{0}}{2}\left[(1+\eta)^{-1/2}+(1-\eta)^{-1/2}\right]\,.$ (7)
The energy shift, $\Delta E_{\mathrm{sc}}=E_{\mathrm{sc}}-\hbar\omega_{0}$, is greater than zero for any value of $\eta$, indicating a repulsive interaction. For $R\gg l$ (where $l$ is the side length of a loop), the value of $\eta$ is much smaller than 1, and $\Delta E_{\mathrm{sc}}$ approximates to
$\Delta E_{\mathrm{sc}}\simeq 3\hbar\omega_{0}\eta^{2}/8.$ (8)
This contrasts with the standard vdW-London interaction, which is caused by capacitive (electric dipole) coupling. In the circuit model, the standard vdW force can be described by a capacitive interaction term Kleppner (1994),
$\frac{\kappa}{C}Q_{1}Q_{2}\,$,
instead of the inductive interaction. The dimensionless coupling constant, $\kappa$, is related to the dipole interaction. For the parallel geometry of the two capacitors, $\kappa=v_{c}/4\pi R^{3}$ at large distances, where $v_{c}$ is the volume of each capacitor. Using the charge variables of Eq. (4a), we can find the two eigenfrequencies: $\omega_{\pm}=\omega_{0}\sqrt{1\pm\kappa}$. Therefore, for a small coupling constant ($\kappa\ll 1$), the change in the zero-point energy is given by $\hbar\omega_{0}\kappa^{2}/8$. This gives rise to the well-known attractive vdW-London interaction, which is proportional to $1/R^{6}$.
Why does the inductive van der Waals-London interaction manifest as a repulsion? Phenomenologically, it can be explained by current correlation. From $I_{1}=(I_{+}+I_{-})/\sqrt{2}$ and $I_{2}=(I_{+}-I_{-})/\sqrt{2}$, the ground-state current correlation is given by
$\langle I_{1}I_{2}\rangle$ $=$ $\frac{1}{2}\left(\langle I_{+}^{2}\rangle-\langle I_{-}^{2}\rangle\right)=\frac{1}{4}\left(\frac{\hbar\omega_{+}}{L+M}-\frac{\hbar\omega_{-}}{L-M}\right)$ (9)
$=$ $\frac{\hbar\omega_{0}}{4L}\left[(1+\eta)^{-3/2}-(1-\eta)^{-3/2}\right],$
which is negative. In other words, the two currents are anticorrelated. Since wires carrying opposite currents repel each other, the interaction is repulsive.
Quantum electrodynamic approach-. In the quantum electrodynamic (QED) approach, the interaction between two loops is indirect. It is mediated by virtual photons (see Fig. 1(b)). The Lagrangian is given by
${\cal L}$ $=$ ${\cal L}_{0}+{\cal L}_{i},$ (10)
${\cal L}_{0}$ $=$ $\frac{L}{2}\dot{Q}_{1}^{2}-V_{1}(Q_{1})+\frac{L}{2}\dot{Q}_{2}^{2}-V_{2}(Q_{2})+{\cal L}_{\mathrm{em}},$
${\cal L}_{i}$ $=$ $\dot{Q}_{1}\oint d\mathbf{r}_{1}\cdot\mathbf{A}+\dot{Q}_{2}\oint d\mathbf{r}_{2}\cdot\mathbf{A}\,.$
Here, ${\cal L}_{0}$ includes the kinetic and the potential energies ($V_{1,2}$) of each loop, as well as the electromagnetic field of the vacuum (${\cal L}_{\mathrm{em}}$). The current in each loop interacts with the radiation gauge field, $\mathbf{A}$. The Hamiltonian, derived via the Legendre transformation and quantization, can be written as follows:
$H$ $=$ $\frac{1}{2L}\left(P_{1}-\oint_{1}\mathbf{A}\cdot d\mathbf{x}\right)^{2}+V_{1}(Q_{1})$ (11)
$+$ $\frac{1}{2L}\left(P_{2}-\oint_{2}\mathbf{A}\cdot d\mathbf{x}\right)^{2}+V_{2}(Q_{2})+H_{\mathrm{em}},$
where
$P_{n}=-i\hbar\frac{\partial}{\partial Q_{n}}$ (12)
is the canonical momentum conjugate to $Q_{n}$ ($n=1,2$), and $H_{\mathrm{em}}$ represents the vacuum electromagnetic field. The vector potential
$\mathbf{A}(\mathbf{x},t)=\sum_{\mathbf{k},\lambda}\alpha_{\mathbf{k}}\left[u_{\mathbf{k}}(\mathbf{x})a_{\mathbf{k}\lambda}e^{-i\omega t}+u_{\mathbf{k}}^{*}(\mathbf{x})a_{\mathbf{k}\lambda}^{\dagger}e^{i\omega t}\right]\hat{e}_{\lambda}\,$, (13)
is expanded in plane wave modes $u_{\mathbf{k}}(\mathbf{x})=e^{i\mathbf{k}\cdot\mathbf{x}}/\sqrt{V}$ with the normalization coefficient of $\alpha_{\mathbf{k}}=\sqrt{\hbar/2\epsilon_{0}\omega}$ and the corresponding angular velocity $\omega=ck$. In the Coulomb gauge, the polarization vector, $\hat{e}_{\lambda}$, is constrained to the transverse modes by the condition $\mathbf{k}\cdot\hat{e}_{\lambda}=0$. The Hamiltonian can be split as $H=H_{0}+H_{\mathrm{in}}$, where
$H_{0}=\frac{1}{2L}P_{1}^{2}+V_{1}(Q_{1})+\frac{1}{2L}P_{2}^{2}+V_{2}(Q_{2})+H_{\mathrm{em}},$ (14a)
is the noninteracting part, and the interaction, $H_{\mathrm{in}}=W+X$, consists of:
$W$ $=$ $-\frac{1}{L}\oint_{1}\mathbf{A}\cdot\mathbf{P}_{1}d\mathbf{x}-\frac{1}{L}\oint_{2}\mathbf{A}\cdot\mathbf{P}_{2}d\mathbf{x},$ (14b)
$X$ $=$ $\frac{1}{2L}\left(\oint_{1}\mathbf{A}\cdot d\mathbf{x}\right)^{2}+\frac{1}{2L}\left(\oint_{2}\mathbf{A}\cdot d\mathbf{x}\right)^{2}.$ (14c)
A key difference from the electric interaction is the presence of the two-photon process, $X\propto\mathbf{A}^{2}$, alongside the single-photon term, $W\propto\mathbf{p}\cdot\mathbf{A}$. Diagrammatic representations are shown in Fig. 2(a). The single-photon process of $W$ can alter the loop state, whereas $X$ does not. The leading-order interaction for states with current flow is the single-photon exchange via $W$ (Fig. 2(b)), which is equivalent to the inductive (magnetic) interaction in classical electrodynamics.
Figure 2: (a) Representation of the interaction between the loop state (solid lines) and the photons (wavy lines). The two types of interaction, $W$ and $X$, involve the single (single wavy line) and double (double wavy line) photons, respectively. (b) Leading-order single-photon exchange through $W$ leads to the magnetic interaction between two current-flowing states.
We calculate the interaction energy in the current-free ground state. The inductive interaction shown in Fig. 2(b) is absent, and the leading-order contributions involve two-photon exchange. Due to two types of perturbation, $W$ and $X$, various terms appear. These terms are depicted schematically in Fig. 3, and the outline of the results from these diagrams is as follows: Here $\Delta E_{\alpha}$ implies the interaction energy calculated from the process $(\alpha)$. Process (i) leads to the standard attractive vdW-London interaction, which is proportional to $1/R^{6}$ when $R$ is much larger than $2a$, where $2a$ is the diameter of the loop. The term $\Delta E_{\mathrm{i-a}}$ is related to the retardation effect, also known as the Casimir-Polder interaction Casimir and Polder (1948). However, this term can be ignored in our system (see the Discussion section). The terms $\Delta E_{\mathrm{ii}}$ and $\Delta E_{\mathrm{iii}}$ are unique to the inductive interaction because they include the two-photon interaction of $X$, which is absent in the electric dipole interaction.
$\Delta E_{\mathrm{i}}$ consists of a kind of fourth-order interactions in $W$ (see Fig. 3), and is given by
$\Delta E_{\mathrm{i}}=\sum_{a,b,c}\frac{\langle 0|W|c\rangle\langle c|W|b\rangle\langle b|W|a\rangle\langle a|W|0\rangle}{(E_{0}-E_{c})(E_{0}-E_{b})(E_{0}-E_{a})}.$ (15)
Here $|a\rangle$, $|b\rangle$, and $|c\rangle$ represent the intermediate states in each step of the process. The energy differences are $E_{0}-E_{a}=-\Delta-\hbar\omega$, $E_{0}-E_{b}=-2\Delta$, and $E_{0}-E_{c}=-\Delta-\hbar\omega^{\prime}$, respectively. Eq. (15) includes self-interaction contributions, which are irrelevant and discarded. We need to calculate two integrals of the form:
$\frac{1}{V}\oint\oint\sum_{\mathbf{k}}\frac{\alpha_{\mathbf{k}}^{2}}{\hbar\omega+\Delta}e^{i\mathbf{k}\cdot(\mathbf{r}_{1}-\mathbf{r}_{2})}d\mathbf{r}_{1}\cdot d\mathbf{r}_{2}.$ (16)
This integral is dominated by the contribution with $\omega\lesssim 2\pi c/|\mathbf{r}_{1}-\mathbf{r}_{2}|$. At short distances ($R\ll 2\pi\hbar c/\Delta$), the denominator can be approximated as $\hbar\omega+\Delta\simeq\hbar\omega$, which allows us to evaluate the integral analytically. This corresponds to the quasi-instantaneous interaction limit (see the Discussion section). Then, we obtain
$\Delta E_{\mathrm{i}}=-\frac{M^{2}}{2\Delta}|\bar{I}_{1}|^{2}|\bar{I}_{2}|^{2},$ (17)
where $\bar{I}_{n}\equiv\langle e|P_{n}|g\rangle/L$ is the matrix element associated with the current at site $n$ ($n=1,2$).
Figure 3: These diagrams illustrate the inductive van der Waals-London interaction between two loops in the current-free ground state via the exchange of two photons. The sum of diagrams (i) and (ii) is equivalent to the semiclassical result. The second-order $X$ process (diagram (iii)) is the predominant contribution. Diagram (i-a), related to the retardation effect, is negligible. Symbols $g$ and $e$ represent the ground and the excited states of a loop, respectively.
Process (ii) is peculiar. It combines the single-photon ($W$) and two-photon ($X$) processes. The interaction energy from this process is given by
$\Delta E_{\mathrm{ii}}=\sum_{a,b}\frac{\langle 0|X|b\rangle\langle b|W|a\rangle\langle a|W|0\rangle+\mathrm{h.c.}}{(E_{0}-E_{b})(E_{0}-E_{a})},$ (18)
where $E_{0}-E_{a}=-(\Delta+\hbar\omega)$ and $E_{0}-E_{b}=-(\hbar\omega+\hbar\omega^{\prime})$. Similarly to $\Delta E_{\mathrm{i}}$, $\Delta E_{\mathrm{ii}}$ can be calculated using the quasi-instantaneous interaction approximation, $\hbar\omega+\Delta\simeq\hbar\omega$, in the denominator of Eq. (18). Additionally, Eq. (18) includes an integral of the form
$\frac{1}{V}\oint\oint\sum_{\mathbf{k}^{\prime}}\frac{\alpha_{\mathbf{k}^{\prime}}^{2}}{\hbar\omega+\hbar\omega^{\prime}}e^{i\mathbf{k}^{\prime}\cdot(\mathbf{r}_{1}-\mathbf{r}_{2})}d\mathbf{r}_{1}\cdot d\mathbf{r}_{2}.$ (19)
The contributions with $\omega^{\prime}\lesssim 2\pi c/|\mathbf{r}_{1}-\mathbf{r}_{2}|$ dominate this integral, so we can neglect $\hbar\omega$ in the denominator. Then, we obtain
$\Delta E_{\mathrm{ii}}=\frac{L}{2}(|\bar{I}_{1}|^{2}+|\bar{I}_{2}|^{2})\eta^{2}.$ (20)
Of the three major contributions, this is the only positive one. It originates from the third-order perturbation of the ground state. Note that the product of two negative energies, $E_{0}-E_{a}$ and $E_{0}-E_{b}$, appears in the denominator of Eq. (18), making the overall sign positive.
Process (iii) is of second order in $X$ and is given by
$\Delta E_{\mathrm{iii}}=\sum_{\gamma}\frac{\langle 0|X|\gamma\rangle\langle\gamma|X|0\rangle+\mathrm{h.c.}}{E_{0}-E_{\gamma}},$ (21)
where $\gamma$ denotes a state with two photons created from the ground state, $|\gamma\rangle=a_{\mathbf{k}\lambda}^{\dagger}a_{\mathbf{k}^{\prime}\lambda^{\prime}}^{\dagger}|0\rangle$. As in the other cases, the self-energy contribution is discarded. The integral of Eq. (19) appears in this term. We use the same approximation as in the calculation of $\Delta E_{\mathrm{ii}}$, $\hbar\omega+\hbar\omega^{\prime}\rightarrow\hbar\omega^{\prime}$, and obtain
$\Delta E_{\mathrm{iii}}=-\frac{\hbar c}{\pi a}\eta(\mathbf{R})\lambda(\mathbf{R}),$ (22a)
where the two dimensionless functions are given by $\eta=M/L$ and
$\lambda(\mathbf{R})\equiv\frac{\mu_{0}a}{4\pi L}\oint_{1}\oint_{2}\frac{d\mathbf{r}_{1}\cdot d\mathbf{r}_{2}}{|\mathbf{r}_{1}-\mathbf{r}_{2}|^{2}},$ (22b)
a dimensionless function depending on the geometry of the system.
Interaction energy and force-. Now, we will analyze the physical implications of these results. $\Delta E_{\mathrm{i}}+\Delta E_{\mathrm{ii}}$ reproduces the semiclassical repulsion. In the circuit model of the loop state in the QED approach, we find that $\Delta=\hbar\omega_{0}$ and $|\bar{I}_{n}|^{2}=\hbar\omega_{0}/2L$. Therefore, we obtain $\Delta E_{\mathrm{i}}=-\hbar\omega_{0}\eta^{2}/8$ and $\Delta E_{\mathrm{ii}}=\hbar\omega_{0}\eta^{2}/2$. This leads to
$\Delta E_{\mathrm{i}}+\Delta E_{\mathrm{ii}}=\frac{3}{8}\hbar\omega_{0}\eta^{2},$ (23)
which coincides with the small-$\eta$ limit of the semiclassical result (8). This explains the origin of the repulsive interaction found in the semiclassical approach. The repulsive interaction is primarily due to the greater value of $\Delta E_{\mathrm{ii}}$ compared to the standard attractive vdW-London interaction energy, $|\Delta E_{\mathrm{i}}|$.
This is not the end of the story. QED contains an additional term, $\Delta E_{\mathrm{iii}}$, which the semiclassical treatment misses. This state-independent term cannot be detected via spectroscopic measurements because the energy shifts are identical for all loop states. They can only be observed via force measurements. In our system, force from this contribution,
$F_{3}\equiv-\partial(\Delta E_{\mathrm{iii}})/\partial R,$ (24)
dominates. In other words,
$|F_{3}/F_{12}|\sim c/(R\omega_{0})\gg 1,$ (25)
in the quasi-instantaneous interaction limit ($R\ll 2\pi\hbar c/\Delta$). Here, $F_{12}\equiv-\partial(\Delta E_{\mathrm{i}}+\Delta E_{\mathrm{ii}})/\partial R$ is the force obtained from the contributions (i) and (ii).
Fig. 4 shows the plots of (a) the semiclassical ground state energy ($\Delta E_{\mathrm{sc}}$; Eq. (7)) and (b) the mutual force ($F_{3}$; Eq. (24)) calculated from the QED approach, both of which are plotted as a function of the distance between the loops. The semiclassical energy leads to a repulsive interaction, whereas the QED result is predominantly attractive. Both results are meaningful in real experiments. The ground state energy is plotted using the semiclassical result, which is equivalent to the QED result of $\Delta E_{\mathrm{i}}+\Delta E_{\mathrm{ii}}$ in the $\eta\ll 1$ limit. This result exhibits a repulsive interaction, which is relevant to measurements involving transitions between the different energy levels. In other words, spectroscopic measurements would reveal this repulsive interaction. Conversely, the mutual force is governed by $F_{3}$, a purely a QED result that shows attractive interactions. This term shifts the energy of all levels equally and cannot therefore be detected by a spectroscopic measurement.
Figure 4: (a) The semiclassical interaction energy, $\Delta E_{\mathrm{sc}}$ (see Eq. (7)), and (b) the mutual force, $F_{3}$ (see Eq. (24)), calculated from the QED approach, as a function of the distance $R$ between the two circular loops with $b/a=0.05$. Note that the dominant QED term yields an attractive mutual force, $F_{3}$, which contrasts sharply with the repulsive interaction energy $\Delta E_{\mathrm{sc}}$, or $\Delta E_{i}+\Delta E_{ii}$ (see Eq. (23)).
Discussions-. The interaction energies exhibit different behaviors at the long ($R\gg 2a$) and short ($R\ll 2a$) distances. For $R\gg 2a$, we find that the semiclassical interaction energy is given by $\Delta E_{\mathrm{sc}}\sim 1/R^{6}$, which is the typical length dependence of the vdW interaction (except that $\Delta E_{\mathrm{sc}}$ leads to a repulsive interaction). However, the QED-specific contribution shows a different behavior: $\Delta E_{\mathrm{iii}}\sim-1/R^{7}$. The short-distance limit is more interesting because (1) it is more accessible to real experiments, and that (2) it cannot be achieved in real atoms. For $R\ll 2a$, we find that
$\Delta E_{\mathrm{sc}}$ $\sim$ $\hbar\omega_{0}\left[\ln{(2a/R)}+\ln{4}-2\right]^{2}$
$\Delta E_{\mathrm{iii}}$ $\sim$ $-\frac{\hbar c}{R}[\ln{(2a/R)}+\ln{4}-2].$
The mutual force, dominated by $F_{3}$, exhibits universal behavior that depends only on $\hbar c/R^{2}$ multiplied by a geometric factor. For $R\ll 2a$, we find that
$F_{3}\simeq-\frac{\hbar c}{R^{2}}g(R),$ (26)
where $g(R)=[\ln{(2a/R)}+\ln{4}-1]/[\ln{(2a/b)}+\ln{4}-2]^{2}$ is a geometry-dependent factor of ${\cal O}(1)$. Here, $b$ denotes the the cross-sectional radius of the loop wire. For a typical mesoscopic scale of about $R=0.1\mathrm{\mu m}$, the magnitude of the force ($\sim\hbar c/R^{2}$) is approximately $3\times 10^{-12}\mathrm{N}$. Current technology can detect this minute force (see, for example, Ref. Moser et al., 2013).
This expression of the force (Eq. (26)) resembles the Casimir force, except for the geometry-specific logarithmic factor of $g(R)$. In one dimension, in particular, the Casimir force is equal to $\hbar c/(24R^{2})$. This suggests a close relationship between the Casimir force and the inductive van der Waals interaction, specifically process (iii) in Fig. 3. It is important to note that the Casimir effect can be derived without reference to zero-point energies, i.e., in terms of quantum forces between charges and currents Schwinger (1975); Jaffe (2005).
We neglected the term $\Delta E_{\mathrm{i-a}}$ (Fig. 3), which is associated with the retardation effect, also known as the Casimir–Polder interaction. This contribution is comparable to $\Delta E_{\mathrm{i}}$ and $\Delta E_{\mathrm{ii}}$ only at large distances, with $R\sim R_{c}=hc/\Delta$ (see, for example, Sec. II.F.2 of Ref. Cohen-Tannoudji et al., 2024). For a typical mesoscopic loop, $\Delta\sim 10^{-5}\mathrm{eV}$, giving $R_{c}\sim 0.1\mathrm{m}$. At this distance, the typical energy scale is $\Delta E_{\mathrm{i-a}}\sim\Delta\eta^{2}={\cal O}(10^{-22})\mathrm{eV}$. This is far below measurable levels. Therefore real experiments could operate in the quasi-instantaneous interaction limit, where this term is negligible.
Conclusion-. In summary, we have investigated the van der Waals-London interaction through inductive coupling between two superconducting quantum loops. Our study reveals a fundamental distinction between the semiclassical and quantum electrodynamic treatments of this mesoscopic system. The semiclassical approach, which is based on the shift in the zero-point energy coupled LC oscillators, predicts a repulsive interaction due to anticorrelated currents. In contrast, the QED treatment shows that the dominant contribution to the mutual force comes from the state-independent two-photon exchange term, $\Delta E_{iii}$, resulting in a predominantly attractive force with a short-distance dependence $F_{3}\sim-\hbar c/R^{2}g(R)$, reminiscent of the Casimir effect. Crucially, the key QED term is invisible in spectroscopy yet governs the mutual force ($|F_{3}/F_{12}|\gg 1$). Our findings reveal the role of quantum fluctuations in the inductive van der Waals force, specifically the $\mathbf{A}^{2}$ term in the Hamiltonian. Our work suggests new experimental avenues for observing the van der Waals force in artificial atoms via controlled mesoscopic circuits.

### v3

@@
 Abstract
-We study the van der Waals-London force, which is typically associated with fluctuating dipoles in atoms, in a mesoscopic circuit consisting of two inductively coupled superconducting loops.
+We study the van der Waals-London force, which is typically associated with fluctuating electric dipoles in atoms, in a mesoscopic circuit consisting of two inductively coupled superconducting loops.
 We investigate the inductive van der Waals-London interaction using both semiclassical and quantum electrodynamic (QED) approaches.
@@
 where $g(R)=[\ln{(2a/R)}+\ln{4}-1]/[\ln{(2a/b)}+\ln{4}-2]^{2}$ is a geometry-dependent factor of ${\cal O}(1)$.
-Here, $b$ denotes the the cross-sectional radius of the loop wire.
+Here, $b$ denotes the cross-sectional radius of the loop wire.
 For a typical mesoscopic scale of about $R=0.1\mathrm{\mu m}$, the magnitude of the force ($\sim\hbar c/R^{2}$) is approximately $3\times 10^{-12}\mathrm{N}$.
@@
 Our study reveals a fundamental distinction between the semiclassical and quantum electrodynamic treatments of this mesoscopic system.
-The semiclassical approach, which is based on the shift in the zero-point energy coupled LC oscillators, predicts a repulsive interaction due to anticorrelated currents.
-In contrast, the QED treatment shows that the dominant contribution to the mutual force comes from the state-independent two-photon exchange term, $\Delta E_{iii}$, resulting in a predominantly attractive force with a short-distance dependence $F_{3}\sim-\hbar c/R^{2}g(R)$, reminiscent of the Casimir effect.
-Crucially, the key QED term is invisible in spectroscopy yet governs the mutual force ($|F_{3}/F_{12}|\gg 1$).
-Our findings reveal the role of quantum fluctuations in the inductive van der Waals force, specifically the $\mathbf{A}^{2}$ term in the Hamiltonian.
+The semiclassical approach, based on the shift in the zero-point energy in coupled LC oscillators, predicts a repulsive interaction due to anticorrelated currents.
+In contrast, the QED treatment shows the opposite result.
+The dominant contribution to the mutual force comes from the state-independent two-photon exchange term, $\Delta E_{iii}$, resulting in a predominantly attractive force with a short-distance dependence $F_{3}\sim-\hbar c/R^{2}g(R)$, reminiscent of the Casimir effect.
+The key QED term is crucial yet invisible in spectroscopy, governing the mutual force ($|F_{3}/F_{12}|\gg 1$).
+The role of vacuum fluctuations in QED has been highlighted through various interesting phenomena such as the Lamb shift, spontaneous emission, and others.
+Our finding is another interesting manifestation of this and reveals the role of quantum fluctuations in the inductive van der Waals force, specifically the $\mathbf{A}^{2}$ term in the Hamiltonian.
+The QED-specific term has a striking effect on the interaction, turning the repulsive force into an attractive one.
 Our work suggests new experimental avenues for observing the van der Waals force in artificial atoms via controlled mesoscopic circuits.
+Furthermore, this could be a new kind of experimental test for the validity of QED.